# Hierarchical Reinforcement Learning for Temporal Abstraction of Listwise Recommendation

### v1

@@ 1 Introduction
In recent years, hybrid Recommender systems (RS) have become increasingly popular to overcome information redundancy and have been widely utilized in a variety of domains (e.g. products, articles, advertisements, music, and movies)  Anidorifón et al. (2015). Hybrid RS selects a list of contents (usually called ‘items’) from overwhelmed candidates for exhibition to meet user preferences, with expected clicks, dwell times or purchases increased. Although with substantial achievement, more advanced recommendation techniques are always necessary because of the complexity of business scenarios, including listwise, cold-start, heterogeneous contents, spatiotemporal-aware (STA) recommendations, etc. Among these, the listwise recommendation is always an important issue, not only due to the wide application of Top-K recommendation  Chen et al. (2019) or feed recommendation  Wu et al. (2021), but also the computational complexity with consideration of mutual-item influence, and position bias.
@@
    Define the user perception/preference as state and the entire recommendation list as action, and model the user interest shift as state transition  Zheng et al. (2018); Chen et al. (2018, 2019); Ie et al. (2019).
@@
    Formulate the listwise ranking problem as a sequential decision-making process, with the state represented by the current list ranked item, while the action is to select the next optimal item to the list Gong et al. (2019); Zhao et al. (2019, 2018).
@@
Here we aim to overcome the above obstacles through better temporal abstractions, by designing a fast-slow learning paradigm Madan et al. (2021) of hierarchical reinforcement learning (HRL). Although there are some previous HRL-based RS efforts Takanobu et al. (2019); Zhao et al. (2020); Xie et al. (2021), they might focus on heterogeneous or multi-goal recommendation problems. Instead, we design our HRL such that one level encodes the user perception and outer-session context, while another level encodes the spatiotemporal status, intra-session context, and mutual-item influences. Besides that, we further improve the sample efficiency by applying Edge Computing or Edge-based AI Zhou et al. (2019b) on our RS. With the rapid development of Edge Computing or Edge AI Zhou et al. (2019b), it is possible to utilize user features and feedback on mobile Gong et al. (2020) or collaborative training with cloud Yao et al. (2021); Chen et al. (2021) in recommendation. We embrace this benefit to further improve the modeling depth of user states and enhance the Markov assumption. We argue that the low-level HRL could be deployed on mobile devices, therefore the on-device features can be involved, training is decoupled, and the communication frequency of cloud service is reduced during model inference. Although this intuition belongs to the paradigm of collaborative training with mobile and cloud Yao et al. (2021); Chen et al. (2021), however, it is the first attempt that tries to train part of RL components on the edge side, to the best of our knowledge Yao et al. (2022).
@@ 2.1 RL-based Recommendation
RL has been widely used in recommendation tasks, while their definitions of MDP can be diversified into two categories. The first categories models the user preference as state and recommendation the entire item list as action, such as Zheng et al. (2018); Chen et al. (2018, 2019); Ie et al. (2019), in which the state transit depicts the user interest shift. On the other hand, the second category is to model the ranking steps of items as state, and the selection of the next favorite item is action. Such methods include MDPrank Xia et al. (2017); Wei et al. (2017), Exact-K Gong et al. (2019), LIRD Zhao et al. (2019) and DeepPage Zhao et al. (2018), which aims to model the mutual influence between items and listwise bias. The intuition of combining the advantages of two types of MDPs motivates our idea of mccHRL.
There are also RL works aiming to fix some special recommendation issues, such as location or POI-based information Zhou et al. (2019a), cold start Wang et al. (2020), heterogeneous items Takanobu et al. (2019); Xie et al. (2021), user long-term engagement Zou et al. (2019) or fairness of items Ge et al. (2021). Our work shares similar interests to the POI-based recommendation with Lian et al. (2020); Zhou et al. (2019a); Luo et al. (2021). However, most of them work on the next-location recommendation given sparse spatial information, to learn a reasonable user-location relationship matrix, while we aim to provide an end-to-end solution of session recommendation with spatial information implicitly considered in our recommending policy.
@@ 2.2 Hierarchical Reinforcement Learning
Hierarchical Reinforcement Learning (HRL) has a hierarchical structure of RL layers in order to solve more complex tasks, reduce the searching space, provide different levels of temporal abstractions, or deal with sparse rewards. HRL methods can be classified into two categories by the coordinated way between different levels. The first is the goal-conditional framework Kulkarni et al. (2016); Vezhnevets et al. (2017); Nachum et al. (2018); Levy et al. (2019), in which a high-level agent learns the goal to drive the low-level agent; the second is sub-task discovery framework, the high-level agent might provide some options (or skill) to reduce the search space of low-level agent Bacon et al. (2017); Florensa et al. (2017). Our methodology belongs to the first category in which the high-level agent studies the user perception embedding as the low-level agent decision basis.
HRL methods also vary in hierarchical structures. Choices include multi-layer DQN Kulkarni et al. (2016), multi-layer policy Nachum et al. (2018), or multi-layer actor-critic Levy et al. (2019). Our model is similar to hierarchical actor-critic (HAC) but reduces its complexity for practical edge deployment.
@@ 2.3 HRL-based Recommendation
There are also attempts to apply Hierarchical Reinforcement Learning (HRL) on recommendation or search Takanobu et al. (2019); Zhao et al. (2020); Xie et al. (2021). Takanobu et al. (2019) uses the High-Level RL as the heterogeneous search source selector; similarly, High-Level RL in HRL-rec Xie et al. (2021) is the content channel selector. MaHRL Zhao et al. (2020) uses the High-Level RL to model the user perception state, to optimize multiple recommendation goals. All of these methods define the Low-Level RL as the item selector, which is similar to our framework. We have a similar definition of High-Level RL with MaHRL, however, our difference includes that (1) our target is to improve a session-based, STA recommendation CTR performance instead of multiple labels; (2) we provide a natural formulation of temporal abstraction to solve the sparse reward issue of listwise recommendation; (3) we deploy the Low-Level part of HRL on the edge side to further improve the methodology throughput.
@@ 2.4 Edge-based Recommendation
Edge Computing (might also be named Edge Intelligence, Edge AI, or on-device AI), in contrast to Cloud computing, has been widely studied recently years Zhou et al. (2019b). Efforts have been made on applications including IoT, 5G, and auto-driving. However, this field still is at its early stage with most efforts focusing on the lightweight and deployment of edge models Yao et al. (2022). For example, EdgeRec Gong et al. (2020) works on the split-deployment which places the memory-consuming embedding module on the cloud while the lightweight recommender is inferenced on the device.
There has been increasing attention on Edge-Cloud collaboration, either privacy-primary such as Federated Learning, or efficient-primary. We are interested in efficient primary methods, to improve the recommender’s personalization. Such efforts including COLLA Lu et al. (2019), DCCL Yao et al. (2021) and MC2-SF Chen et al. (2021). For example, COLLA designed the cloud model as a global aggregator distilling knowledge from many edge models; MC2-SF proposes a slow-fast learning mechanism on RS, with the bidirectional collaboration between Edge and Cloud. However, edge-based RL has not been widely studied so far. Instead, RL is often utilized as an aside system to Edge Computing, to help service offloading, task scheduling, or resource allocation Yao et al. (2022). Therefore, our work can be regarded as the pioneering study on edge-based RL, with the slow-fast collaborative study mechanism between cloud and edge, similar to Madan et al. (2021) and Chen et al. (2021). To differentiate from the traditional edge AI methods, we use the notation ‘mobile’ instead of ‘edge’ since most impressions of internet RS are on the users’ mobile devices (smartphones or pads).
@@ 3.2 Reinforcement Learning
RL optimizes a long-term objective which is defined as the discounted accumulated rewards $J=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma\in[0,1)$ is the discount factor. The goal of reinforcement learning is to learn a policy $\pi(a|s)$ which maximizes $J$. In this work, we employ the classical model-free and off-policy algorithm called the Deep Deterministic Policy Gradient (DDPG) method Lillicrap et al. (2016) to solve this problem.
@@ 3.3 Deep Deterministic Policy Gradient
In this paper, we implement RL with a classical model-free and off-policy algorithm called the Deep Deterministic Policy Gradient (DDPG) method Lillicrap et al. (2016). Here we briefly review its algorithmic details which is the basis of later derivation.
@@ 3.4 Gated Recurrent Unit
Gated Recurrent Unit (GRU) is a special type of RNN that can sequentially capture the item-dependency impacts. It is found that GRU has better stability with fewer parameters than general GRN, therefore is usually used in recommendation studies Hidasi et al. (2016); Zhao et al. (2019); Zhou et al. (2019b). The core logic of GRU is the following:
@@ 3.5 The mccHRL Framework
  • Low-Level Reward $r^{\text{l}}$: the intrinsic motivation received by LRA. Here we embrace the goal-conditioned framework of HRL Vezhnevets et al. (2017); Nachum et al. (2018); Nasiriany et al. (2019), i.e., the nearest action of HRA studies the current user preference and therefore can be used to calculate the goal of LRA action. For instance, the LRA reward $r^{\text{l}}$ can be simply formulated as the Ellucidian distance between the HRA action and LRA action.
@@ 4.2 High-Level Agent
The HRA performs as the critic in our mccHRL. It is modeled with a parametric value function $Q_{w}(s^{\text{h}},a^{\text{h}})$. The left part of Figure 3 shows the critic network structure. Historical click sequences $\{e\}$ are extracted from the cloud database, each encoded by SRE and then processed by a DIN structure Zhou et al. (2018) with the last $l$ as the target tensor. The DIN output concatenates with $u$ and $c^{o}$ and goes into first an FM layer first then an MLP. The last activation function of MLP is tanh which generates the Q-value. At the beginning of a session service, HRA also transmits $u$, $c^{o}$, and SRE to LRA, as indicated by Figure 3. The output of $Q_{w}$ then provides an implicit function of LRA goals $g$, as indicated by Arrow III.
@@ 4.4.2 Training with Offline Data
RL study on such data calls for a special research direction called offline RL, or batch-constrained RL, which learns an optimal policy from a fixed dataset by implementing the out-of-distribution corrections for off-policy RL Levine et al. (2020). Since our HRA has an off-policy strategy, we could employ similar ideas to train the HRA critic offline. Here we refer to the conservative Q-learning (CQL) Kumar et al. (2020), to have a regularizer in the TD loss of the Q-function:
@@
in which $Z$ is the normalizing factor of $Q$ and $\pi_{\beta}$ is the sampling policy in data. This formulation indicates that we also minimize the soft-maximum of Q at the current solved policy ($\pi_{\theta}(a|s)$) w.r.t the sampling policy $a^{\text{h}}$, therefore the learning becomes ’conservative’. According to Algorithm 1 in Kumar et al. (2020), such a trick can be also applied in the actor-critic framework. Here we present our offline version of Algorithm 1 which is specified in Algorithm 2.
@@ 5.1 Experimental Environments
Based on the MovieLens-100k dataset, We build a simulator based on the methodology introduced in Zhao et al. (2019), with the averaged ratings behaving as the RL rewards. The average rating is 3.53 over all samples. Among the user features, the spatial features including ‘occupation’ and ‘zip code’ are set as edge features, which can not be accessed by the cloud. We further assume the user cloud sequence has a fixed latency with edge sequence, i.e., the cloud sequence is delayed by 6 items the edge sequence. By these configurations, we build an approximated mobile-cloud simulator, with an explicit transmission latency.
@@ 5.4 Baselines
  • DIN: the Deep Interest Network Zhou et al. (2018), as a standardized solution of industrial pointwise CTR model.
@@
  • GRU4rec: a session-based method with a latent state calculated by GRU Hidasi et al. (2016), therefore considers the intra-session context.
  • LIRD: a DDPG-based RL listwise ranking framework Zhao et al. (2019).
  • MC2-SF: a slow-fast training framework based on moble-cloud collaboration Chen et al. (2021), with mobile features incorporated and considered.

### v2

@@ 1 Introduction
In recent years, hybrid Recommender systems (RS) have become increasingly popular to overcome information redundancy and have been widely utilized in a variety of domains (e.g. products, articles, advertisements, music, and movies)  (Anidorifón et al., 2015). Hybrid RS selects a list of contents (usually called ‘items’) from overwhelmed candidates for exhibition to meet user preferences, with expected clicks, dwell times or purchases increased. Although with substantial achievement, more advanced recommendation techniques are always necessary because of the complexity of business scenarios, including listwise, cold-start, heterogeneous contents, spatiotemporal-aware (STA) recommendations, etc. Among these, the listwise recommendation is always an important issue, not only due to the wide application of Top-K recommendation  (Chen et al., 2019) or feed recommendation  (Wu et al., 2021), but also the computational complexity with consideration of mutual-item influence, and position bias.
@@
    Define the user perception/preference as state and the entire recommendation list as action, and model the user interest shift as state transition  (Zheng et al., 2018; Chen et al., 2018, 2019; Ie et al., 2019).
@@
    Formulate the listwise ranking problem as a sequential decision-making process, with the state represented by the current list ranked item, while the action is to select the next optimal item to the list (Gong et al., 2019; Zhao et al., 2019, 2018).
@@
Here we aim to overcome the above obstacles through better temporal abstractions, by designing a fast-slow learning paradigm (Madan et al., 2021) of hierarchical reinforcement learning (HRL). Although there are some previous HRL-based RS efforts (Takanobu et al., 2019; Zhao et al., 2020; Xie et al., 2021), they might focus on heterogeneous or multi-goal recommendation problems. Instead, we design our HRL such that one level encodes the user perception and outer-session context, while another level encodes the spatiotemporal status, intra-session context, and mutual-item influences. Besides that, we further improve the sample efficiency by applying Edge Computing or Edge-based AI (Zhou et al., 2019b) on our RS. With the rapid development of Edge Computing or Edge AI (Zhou et al., 2019b), it is possible to utilize user features and feedback on mobile (Gong et al., 2020) or collaborative training with cloud (Yao et al., 2021; Chen et al., 2021) in recommendation. We embrace this benefit to further improve the modeling depth of user states and enhance the Markov assumption. We argue that the low-level HRL could be deployed on mobile devices, therefore the on-device features can be involved, training is decoupled, and the communication frequency of cloud service is reduced during model inference. Although this intuition belongs to the paradigm of collaborative training with mobile and cloud (Yao et al., 2021; Chen et al., 2021), however, it is the first attempt that tries to train part of RL components on the edge side, to the best of our knowledge (Yao et al., 2022).
@@ 2.1 RL-based Recommendation
RL has been widely used in recommendation tasks, while their definitions of MDP can be diversified into two categories. The first categories models the user preference as state and recommendation the entire item list as action, such as (Zheng et al., 2018; Chen et al., 2018, 2019; Ie et al., 2019), in which the state transit depicts the user interest shift. On the other hand, the second category is to model the ranking steps of items as state, and the selection of the next favorite item is action. Such methods include MDPrank (Xia et al., 2017; Wei et al., 2017), Exact-K (Gong et al., 2019), LIRD (Zhao et al., 2019) and DeepPage (Zhao et al., 2018), which aims to model the mutual influence between items and listwise bias. The intuition of combining the advantages of two types of MDPs motivates our idea of mccHRL.
There are also RL works aiming to fix some special recommendation issues, such as location or POI-based information (Zhou et al., 2019a), cold start (Wang et al., 2020), heterogeneous items (Takanobu et al., 2019; Xie et al., 2021), user long-term engagement (Zou et al., 2019) or fairness of items (Ge et al., 2021). Our work shares similar interests to the POI-based recommendation with (Lian et al., 2020; Zhou et al., 2019a; Luo et al., 2021). However, most of them work on the next-location recommendation given sparse spatial information, to learn a reasonable user-location relationship matrix, while we aim to provide an end-to-end solution of session recommendation with spatial information implicitly considered in our recommending policy.
@@ 2.2 Hierarchical Reinforcement Learning
Hierarchical Reinforcement Learning (HRL) has a hierarchical structure of RL layers in order to solve more complex tasks, reduce the searching space, provide different levels of temporal abstractions, or deal with sparse rewards. HRL methods can be classified into two categories by the coordinated way between different levels. The first is the goal-conditional framework (Kulkarni et al., 2016; Vezhnevets et al., 2017; Nachum et al., 2018; Levy et al., 2019), in which a high-level agent learns the goal to drive the low-level agent; the second is sub-task discovery framework, the high-level agent might provide some options (or skill) to reduce the search space of low-level agent (Bacon et al., 2017; Florensa et al., 2017). Our methodology belongs to the first category in which the high-level agent studies the user perception embedding as the low-level agent decision basis.
HRL methods also vary in hierarchical structures. Choices include multi-layer DQN (Kulkarni et al., 2016), multi-layer policy (Nachum et al., 2018), or multi-layer actor-critic (Levy et al., 2019). Our model is similar to hierarchical actor-critic (HAC) but reduces its complexity for practical edge deployment.
@@ 2.3 HRL-based Recommendation
There are also attempts to apply Hierarchical Reinforcement Learning (HRL) on recommendation or search (Takanobu et al., 2019; Zhao et al., 2020; Xie et al., 2021). (Takanobu et al., 2019) uses the High-Level RL as the heterogeneous search source selector; similarly, High-Level RL in HRL-rec (Xie et al., 2021) is the content channel selector. MaHRL (Zhao et al., 2020) uses the High-Level RL to model the user perception state, to optimize multiple recommendation goals. All of these methods define the Low-Level RL as the item selector, which is similar to our framework. We have a similar definition of High-Level RL with MaHRL, however, our difference includes that (1) our target is to improve a session-based, STA recommendation CTR performance instead of multiple labels; (2) we provide a natural formulation of temporal abstraction to solve the sparse reward issue of listwise recommendation; (3) we deploy the Low-Level part of HRL on the edge side to further improve the methodology throughput.
@@ 2.4 Edge-based Recommendation
Edge Computing (might also be named Edge Intelligence, Edge AI, or on-device AI), in contrast to Cloud computing, has been widely studied recently years (Zhou et al., 2019b). Efforts have been made on applications including IoT, 5G, and auto-driving. However, this field still is at its early stage with most efforts focusing on the lightweight and deployment of edge models (Yao et al., 2022). For example, EdgeRec (Gong et al., 2020) works on the split-deployment which places the memory-consuming embedding module on the cloud while the lightweight recommender is inferenced on the device.
There has been increasing attention on Edge-Cloud collaboration, either privacy-primary such as Federated Learning, or efficient-primary. We are interested in efficient primary methods, to improve the recommender’s personalization. Such efforts including COLLA (Lu et al., 2019), DCCL (Yao et al., 2021) and MC2-SF (Chen et al., 2021). For example, COLLA designed the cloud model as a global aggregator distilling knowledge from many edge models; MC2-SF proposes a slow-fast learning mechanism on RS, with the bidirectional collaboration between Edge and Cloud. However, edge-based RL has not been widely studied so far. Instead, RL is often utilized as an aside system to Edge Computing, to help service offloading, task scheduling, or resource allocation (Yao et al., 2022). Therefore, our work can be regarded as the pioneering study on edge-based RL, with the slow-fast collaborative study mechanism between cloud and edge, similar to (Madan et al., 2021) and (Chen et al., 2021). To differentiate from the traditional edge AI methods, we use the notation ‘mobile’ instead of ‘edge’ since most impressions of internet RS are on the users’ mobile devices (smartphones or pads).
@@ 3.2 Reinforcement Learning
RL optimizes a long-term objective which is defined as the discounted accumulated rewards $J=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma\in[0,1)$ is the discount factor. The goal of reinforcement learning is to learn a policy $\pi(a|s)$ which maximizes $J$. In this work, we employ the classical model-free and off-policy algorithm called the Deep Deterministic Policy Gradient (DDPG) method (Lillicrap et al., 2016) to solve this problem.
@@ 3.3 Deep Deterministic Policy Gradient
In this paper, we implement RL with a classical model-free and off-policy algorithm called the Deep Deterministic Policy Gradient (DDPG) method (Lillicrap et al., 2016). Here we briefly review its algorithmic details which is the basis of later derivation.
@@ 3.4 Gated Recurrent Unit
Gated Recurrent Unit (GRU) is a special type of RNN that can sequentially capture the item-dependency impacts. It is found that GRU has better stability with fewer parameters than general GRN, therefore is usually used in recommendation studies (Hidasi et al., 2016; Zhao et al., 2019; Zhou et al., 2019b). The core logic of GRU is the following:
@@ 3.5 The mccHRL Framework
  • Low-Level Reward $r^{\text{l}}$: the intrinsic motivation received by LRA. Here we embrace the goal-conditioned framework of HRL (Vezhnevets et al., 2017; Nachum et al., 2018; Nasiriany et al., 2019), i.e., the nearest action of HRA studies the current user preference and therefore can be used to calculate the goal of LRA action. For instance, the LRA reward $r^{\text{l}}$ can be simply formulated as the Ellucidian distance between the HRA action and LRA action.
@@ 4.2 High-Level Agent
The HRA performs as the critic in our mccHRL. It is modeled with a parametric value function $Q_{w}(s^{\text{h}},a^{\text{h}})$. The left part of Figure 3 shows the critic network structure. Historical click sequences $\{e\}$ are extracted from the cloud database, each encoded by SRE and then processed by a DIN structure (Zhou et al., 2018) with the last $l$ as the target tensor. The DIN output concatenates with $u$ and $c^{o}$ and goes into first an FM layer first then an MLP. The last activation function of MLP is tanh which generates the Q-value. At the beginning of a session service, HRA also transmits $u$, $c^{o}$, and SRE to LRA, as indicated by Figure 3. The output of $Q_{w}$ then provides an implicit function of LRA goals $g$, as indicated by Arrow III.
@@ 4.4.2 Training with Offline Data
RL study on such data calls for a special research direction called offline RL, or batch-constrained RL, which learns an optimal policy from a fixed dataset by implementing the out-of-distribution corrections for off-policy RL (Levine et al., 2020). Since our HRA has an off-policy strategy, we could employ similar ideas to train the HRA critic offline. Here we refer to the conservative Q-learning (CQL) (Kumar et al., 2020), to have a regularizer in the TD loss of the Q-function:
@@
in which $Z$ is the normalizing factor of $Q$ and $\pi_{\beta}$ is the sampling policy in data. This formulation indicates that we also minimize the soft-maximum of Q at the current solved policy ($\pi_{\theta}(a|s)$) w.r.t the sampling policy $a^{\text{h}}$, therefore the learning becomes ’conservative’. According to Algorithm 1 in (Kumar et al., 2020), such a trick can be also applied in the actor-critic framework. Here we present our offline version of Algorithm 1 which is specified in Algorithm 2.
@@ 5.1 Experimental Environments
Based on the MovieLens-100k dataset, We build a simulator based on the methodology introduced in (Zhao et al., 2019), with the averaged ratings behaving as the RL rewards. The average rating is 3.53 over all samples. Among the user features, the spatial features including ‘occupation’ and ‘zip code’ are set as edge features, which can not be accessed by the cloud. We further assume the user cloud sequence has a fixed latency with edge sequence, i.e., the cloud sequence is delayed by 6 items the edge sequence. By these configurations, we build an approximated mobile-cloud simulator, with an explicit transmission latency.
@@ 5.4 Baselines
  • DIN: the Deep Interest Network (Zhou et al., 2018), as a standardized solution of industrial pointwise CTR model.
@@
  • GRU4rec: a session-based method with a latent state calculated by GRU (Hidasi et al., 2016), therefore considers the intra-session context.
  • LIRD: a DDPG-based RL listwise ranking framework (Zhao et al., 2019).
  • MC2-SF: a slow-fast training framework based on moble-cloud collaboration (Chen et al., 2021), with mobile features incorporated and considered.